# Spectrum of Schrödinger operators on subcovering graphs

### v2

@@ 4. Examples
In this section we consider simple examples of periodic graphs and describe the spectra of the Schrödinger operators on their subcoverings.
@@ 4.1. Nanotubes
The hexagonal lattice ${\bf G}$ is a $\Gamma$-periodic graph, where $\Gamma\subset{\mathbb{R}}^{2}$ is the lattice with the basis $\mathfrak{a}_{1},\mathfrak{a}_{2}$, see Fig. 1a. The fundamental graph ${\bf G}_{*}$ of ${\bf G}$ consists of two vertices $v_{1},v_{2}$ and three multiple edges connecting these vertices, see Fig. 1d.
First we briefly describe the spectrum of the Schrödinger operator $H=\Delta+Q$ on ${\bf G}$ with a $\Gamma$-periodic potential $Q$ defined by (2.7), for more details see, e.g., [KP07]. The Floquet operator $H(k)$ given by (1.4) in the standard basis of $\ell^{2}(\{v_{1},v_{2}\})\cong{\mathbb{C}}^{2}$ has the following matrix:
@@
Proof of Example 2.6. Due to Proposition 2.2, the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on the nanotube ${\bf G}_{\mathfrak{t}}$ with a primitive chiral vector $\mathfrak{t}=t_{1}\mathfrak{a}_{1}+t_{2}\mathfrak{a}_{2}$ consists of two bands:
@@
ii) Let $t_{1}-t_{2}\not\in 3{\mathbb{Z}}$, i.e., $t_{1}-t_{2}=3m\pm 1$ for some $m\in{\mathbb{Z}}$. We obtain the asymptotics of the band edge $\lambda_{1}^{+}(\mathfrak{t})$ as $|t|\to\infty$, where $t=(t_{1},t_{2})$. The proof for $\lambda_{2}^{-}(\mathfrak{t})$ is similar. The point $k_{o}=\big{(}\frac{2\pi}{3}\,,-\frac{2\pi}{3}\big{)}$ is the only (up to evenness) maximum point of the band function $\lambda_{1}(k)$ in ${\mathcal{B}}=(-\pi,\pi]^{2}$, see (4.2), (4.3). Using the identity (4.1) for $\lambda_{1}(k)$, we obtain
@@
Thus, the band function $\lambda_{1}(k)$ of $H$ satisfies the Assumption A and we can use the asymptotics (2.6). For $t=(t_{1},t_{2})$ and $k_{o}=\big{(}\frac{2\pi}{3}\,,-\frac{2\pi}{3}\big{)}$, we obtain
$|{\bf H}^{-1/2}t|^{2}=\langle t,{\bf H}^{-1}t\rangle=\frac{4q}{3}\,(t_{1}^{2}+t_{1}t_{2}+t_{2}^{2}),$ (4.4)
$\langle t,k_{o}\rangle=\frac{2\pi}{3}\,(t_{1}-t_{2}).$
@@
$x_{o}:=(\langle t,k_{o}\rangle\mod 2\pi)=\pm\frac{2\pi}{3}\,.$ (4.5)
@@
$\lambda_{1}^{+}(\mathfrak{t})=3-q-\frac{\pi^{2}}{6q(t_{1}^{2}+t_{1}t_{2}+t_{2}^{2})}+O\big{(}\frac{1}{|t|^{3}}\big{)}\qquad\textrm{as}\qquad|t|\to\infty.\qquad\Box$
@@ 4.2. Subcoverings of the $d$-dimensional lattice
The $d$-dimensional lattice ${\mathbb{L}}^{d}$ is a ${\mathbb{Z}}^{d}$-periodic graph, for $d=3$ see Fig. 2a. The fundamental graph ${\mathbb{L}}^{d}_{*}$ of ${\mathbb{L}}^{d}$ consists of one vertex $v$ and $d$ loop edges at this vertex $v$ (Fig. 2d). The Floquet Laplacian $\Delta(k)$ for the lattice ${\mathbb{L}}^{d}$ has the form
@@
Finally, if $d_{o}=1$ and $\varrho_{1}$ is odd, then ${\mathcal{T}}=\mathfrak{t}_{1}^{\top}$, $\varrho=1$. Substituting this into (2.10), we obtain (2.11).  $\Box$
@@ 4.3. Subcoverings of the diamond lattice
of the hexagonal lattice ${\bf G}\subset{\mathbb{R}}^{2}$ along the vector $\mathfrak{a}_{3}=(0,0,1)$. Here ${\mathbb{R}}^{2}$ is considered as the subspace ${\mathbb{R}}^{2}\times\{0\}$ of ${\mathbb{R}}^{3}$. The copies of ${\bf G}$ are connected in a periodic way by the edges
$\{v_{1},v_{2}+\mathfrak{a}_{3}\}+\mathfrak{a},\qquad\forall\,\mathfrak{a}\in\Gamma,$
between the vertex $v_{1}+\mathfrak{a}$ in a lower copy and the vertex $v_{2}+\mathfrak{a}_{3}+\mathfrak{a}$ in the next copy of ${\bf G}$, $\mathfrak{a}\in\Gamma$, where $v_{1},v_{2}$ are the vertices of ${\bf G}$ from the fundamental cell $\Omega$, and $\Gamma$ is the lattice with the basis $\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}$. The diamond lattice ${\mathcal{G}}$ is a $\Gamma$-periodic graph with the fundamental graph ${\mathcal{G}}_{*}$ consisting of two vertices $v_{1},v_{2}$ and four multiple edges connecting these vertices, see Fig. 3d.
The spectrum of the Schrödinger operator $H=\Delta+Q$ on the diamond lattice ${\mathcal{G}}$ with a $\Gamma$-periodic potential $Q$ defined by (2.12) is given by
@@
Proof of Example 2.10. Let $\mathfrak{t}=t_{1}\mathfrak{a}_{1}+t_{2}\mathfrak{a}_{2}+t_{3}\mathfrak{a}_{3}$ be a primitive vector of the lattice $\Gamma$. For some $i,j=1,2,3$, $i\neq j$, the numbers $t_{i}$ and $t_{j}$ have the same parity, i.e., $t_{i}+t_{j}$ is even. Let $k_{o}=(k_{o1},k_{o2},k_{o3})$ be defined by
@@
Then $k_{o}\in K_{1}^{+}=K_{2}^{-}$, where $K_{1}^{+}=K_{2}^{-}$ is defined by (4.7) and for $k_{o}$ and ${\mathcal{T}}=(t_{1},t_{2},t_{3})$ the condition (2.3) is fulfilled. For the point $0\in K_{1}^{-}=K_{2}^{+}$ the condition (2.3) is also fulfilled. Thus, due to Proposition 2.2.i, the diamond lattice ${\mathcal{G}}$ is isospectral to its subcovering ${\mathcal{G}}_{\mathfrak{t}}$, i.e., $\sigma(H)=\sigma(H_{\mathfrak{t}})$.
Now let $\mathfrak{t}=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$ be a primitive set of the lattice $\Gamma$, where $\mathfrak{t}_{s}=t_{s1}\mathfrak{a}_{1}+t_{s2}\mathfrak{a}_{2}+t_{s3}\mathfrak{a}_{3}$, $s=1,2$. If for some $i,j=1,2,3$, $i\neq j$, the sum $t_{si}+t_{sj}$ is even for each $s=1,2$, then using the same arguments as above we obtain that ${\mathcal{G}}$ and ${\mathcal{G}}_{\mathfrak{t}}$ are isospectral. Conversely, if ${\mathcal{G}}$ and ${\mathcal{G}}_{\mathfrak{t}}$ are isospectral, then, due to Proposition 2.2.i,
@@
Acknowledgments. The work of the author was supported by the RSF grant No. 25-21-00157.

### v3

@@ 4. Examples
In this section we consider simple examples of periodic graphs (with the unit edge weight) and describe the spectra of the Schrödinger operators on their subcoverings.
@@ 4.1. Nanotubes
The hexagonal lattice ${\bf G}$ is a ${\mathbb{Z}}^{2}$-periodic graph with the periods $\mathfrak{a}_{1},\mathfrak{a}_{2}$, see Fig. 1a. The fundamental graph ${\bf G}_{*}$ of ${\bf G}$ consists of two vertices $v_{1},v_{2}$ and three multiple edges connecting these vertices, see Fig. 1d.
First we briefly describe the spectrum of the Schrödinger operator $H=\Delta+Q$ on ${\bf G}$ with a ${\mathbb{Z}}^{2}$-periodic potential $Q$ defined by (2.8), for more details see, e.g., [KP07]. The Floquet operator $H(k)$ given by (1.4) in the standard basis of $\ell^{2}(\{v_{1},v_{2}\})\cong{\mathbb{C}}^{2}$ has the following matrix:
@@
Proof of Example 2.6. Due to Proposition 2.2, the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on the nanotube ${\bf G}_{\mathfrak{t}}$ with a primitive chiral vector $\mathfrak{t}=(t_{1},t_{2})\in{\mathbb{Z}}^{2}$ consists of two bands:
@@
ii) Let $t_{1}-t_{2}\not\in 3{\mathbb{Z}}$, i.e., $t_{1}-t_{2}=3m\pm 1$ for some $m\in{\mathbb{Z}}$. We obtain the asymptotics of the band edge $\lambda_{1}^{+}(\mathfrak{t})$ as $|\mathfrak{t}|\to\infty$. The proof for $\lambda_{2}^{-}(\mathfrak{t})$ is similar. The point $k_{o}=\big{(}\frac{2\pi}{3}\,,-\frac{2\pi}{3}\big{)}$ is the only (up to evenness) maximum point of the band function $\lambda_{1}(k)$ in ${\mathcal{B}}=(-\pi,\pi]^{2}$, see (4.2), (4.3). Using the identity (4.1) for $\lambda_{1}(k)$, we obtain
@@
Thus, the band function $\lambda_{1}(k)$ of $H$ satisfies the Assumption A and we can use the asymptotics (2.7). For $\mathfrak{t}=(t_{1},t_{2})$ and $k_{o}=\big{(}\frac{2\pi}{3}\,,-\frac{2\pi}{3}\big{)}$, we obtain
$|{\bf H}^{-1/2}\mathfrak{t}|^{2}=\langle\mathfrak{t},{\bf H}^{-1}\mathfrak{t}\rangle=\frac{4q}{3}\,(t_{1}^{2}+t_{1}t_{2}+t_{2}^{2}),$ (4.4)
$\langle\mathfrak{t},k_{o}\rangle=\frac{2\pi}{3}\,(t_{1}-t_{2}).$
@@
$x_{o}:=(\langle\mathfrak{t},k_{o}\rangle\mod 2\pi)=\pm\frac{2\pi}{3}\,.$ (4.5)
@@
$\lambda_{1}^{+}(\mathfrak{t})=3-q-\frac{\pi^{2}}{6q(t_{1}^{2}+t_{1}t_{2}+t_{2}^{2})}+O\big{(}\frac{1}{|\mathfrak{t}|^{3}}\big{)}\qquad\textrm{as}\qquad|\mathfrak{t}|\to\infty.\qquad\Box$
@@ 4.2. Subcoverings of the $d$-dimensional lattice
The $d$-dimensional lattice ${\mathbb{L}}^{d}$ is a ${\mathbb{Z}}^{d}$-periodic graph with the periods $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}$, for $d=3$ see Fig. 2a. The fundamental graph ${\mathbb{L}}^{d}_{*}$ of ${\mathbb{L}}^{d}$ consists of one vertex $v$ and $d$ loop edges at this vertex $v$ (Fig. 2d). The Floquet Laplacian $\Delta(k)$ for the lattice ${\mathbb{L}}^{d}$ has the form
@@
Finally, if ${\mathbb{L}}^{d}_{\mathfrak{t}}$ is a subcovering of ${\mathbb{L}}^{d}$ with a single primitive chiral vector $\mathfrak{t}=(t_{1},\ldots,t_{d})\in{\mathbb{Z}}^{d}$ and $t_{1}+\ldots+t_{d}$ is odd, then
${\mathcal{T}}=\mathfrak{t}^{\top},\qquad\varrho=(t_{1}+\ldots+t_{d}\mod 2)=1.$
Substituting this into (2.11), we obtain (2.12).  $\Box$
@@ 4.3. Subcoverings of the diamond lattice
of the hexagonal lattice ${\bf G}\subset{\mathbb{R}}^{2}$ along the vector $\mathfrak{a}_{3}$. Here ${\mathbb{R}}^{2}$ is considered as the subspace ${\mathbb{R}}^{2}\times\{0\}$ of ${\mathbb{R}}^{3}$. The copies of ${\bf G}$ are connected in a periodic way by the edges between white vertices in a lower copy and black vertices in the next copy of ${\bf G}$. The diamond lattice ${\mathcal{G}}$ is a ${\mathbb{Z}}^{3}$-periodic graph with the periods $\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}$, where $\mathfrak{a}_{1},\mathfrak{a}_{2}$ are the periods of the hexagonal lattice ${\bf G}$. The fundamental graph ${\mathcal{G}}_{*}$ of ${\mathcal{G}}$ consists of two vertices $v_{1},v_{2}$ and four multiple edges connecting these vertices, see Fig. 3d.
The spectrum of the Schrödinger operator $H=\Delta+Q$ on the diamond lattice ${\mathcal{G}}$ with a ${\mathbb{Z}}^{3}$-periodic potential $Q$ defined by (2.13) is given by
@@
Proof of Example 2.10. Let $\mathfrak{t}=(t_{1},t_{2},t_{3})\in{\mathbb{Z}}^{3}$ be a primitive vector. For some $i,j=1,2,3$, $i\neq j$, the numbers $t_{i}$ and $t_{j}$ have the same parity, i.e., $t_{i}+t_{j}$ is even. Let $k_{o}=(k_{o1},k_{o2},k_{o3})$ be defined by
@@
Then $k_{o}\in K_{1}^{+}=K_{2}^{-}$, where $K_{1}^{+}=K_{2}^{-}$ is defined by (4.7) and for $k_{o}$ and ${\mathcal{T}}=\mathfrak{t}^{\top}$ the condition (2.3) is fulfilled. For the point $0\in K_{1}^{-}=K_{2}^{+}$ the condition (2.3) is also fulfilled. Thus, due to Proposition 2.2.i, the diamond lattice ${\mathcal{G}}$ is isospectral to its subcovering ${\mathcal{G}}_{\mathfrak{t}}$, i.e., $\sigma(H)=\sigma(H_{\mathfrak{t}})$.
Now let $\mathfrak{t}=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}\subset{\mathbb{Z}}^{3}$ be a primitive set, where $\mathfrak{t}_{s}=(t_{s1},t_{s2},t_{s3})$, $s=1,2$. If for some $i,j=1,2,3$, $i\neq j$, the sum $t_{si}+t_{sj}$ is even for each $s=1,2$, then using the same arguments as above we obtain that ${\mathcal{G}}$ and ${\mathcal{G}}_{\mathfrak{t}}$ are isospectral. Conversely, if ${\mathcal{G}}$ and ${\mathcal{G}}_{\mathfrak{t}}$ are isospectral, then, due to Proposition 2.2.i,
@@
Acknowledgments. This work is supported by the Russian Science Foundation (project No. 25-21-00157). I would like to thank the referees for thoughtful comments that helped me to improve the manuscript.